An electron is doubled from .2c to .4c.By what ratio does it's momentum increease?(b)What happens to the momentum ratio when the electron speed is doubled again from 0.4c to 0.8c.
1
Expert's answer
2019-08-14T09:44:45-0400

a) Relativistic momentum:


"p=\frac{mv}{\sqrt{1-(\frac{v}{c})^2}}"

The ratio will be


"\frac{\frac{0.4}{\sqrt{1-0.4^2}}}{\frac{0.2}{\sqrt{1-0.2^2}}}=2.138"

b) The ratio will be


"\frac{\frac{0.8}{\sqrt{1-0.8^2}}}{\frac{0.4}{\sqrt{1-0.4^2}}}=3.055"